OpenStudy (anonymous):

Which of the following shows a proportion? 4/9 = 18/8 3/8 = 6/4 6/12 =8/4 9/15 = 12/20

OpenStudy (anonymous):

@zimmah

OpenStudy (anonymous):

the last one, the other ones are all reversed

OpenStudy (anonymous):

4/9 = 8/18 3/8 = 4/6 6/12 = 4/8

OpenStudy (anonymous):

or actually no 3/8 is not 4/6

OpenStudy (anonymous):

but also not 6/4

OpenStudy (anonymous):

that one just is an oddity

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

9/15 = 12/20 all the other ones are false. 9/15 = 0.6 12/20 = 0.6 so they are equal all the others are no equal, so they are false

OpenStudy (anonymous):

Thank you very much